OpenStudy (anonymous):

Verify the identity. cot(x-pi/2) = -tanx

OpenStudy (anonymous):

Ok?

OpenStudy (anonymous):

Sorry it didn't let me type it all out. cot(x-pi/2) = -tanx

OpenStudy (jdoe0001):

blank, why? because "blank" = "blank"

OpenStudy (anonymous):

how do I go about proving it?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ cot(\theta)=\cfrac{cos(\theta)}{sin(\theta)}}}\\ \quad \\ \quad \\ cot\left(x-\frac{\pi}{2}\right)=-tan(x)\\ \quad \\ cot\left(x-\frac{\pi}{2}\right)\implies \cfrac{cos\left(x-\frac{\pi}{2}\right)}{sin\left(x-\frac{\pi}{2}\right)}\\ \quad \\ \implies \cfrac{cos(x)cos\left(\frac{\pi}{2}\right)+sin(x)sin\left(\frac{\pi}{2}\right)}{sin(x)cos\left(\frac{\pi}{2}\right)-cos(x)sin\left(\frac{\pi}{2}\right)}\) notice your Unit Circle, what's \(\bf cos\left(\frac{\pi}{2}\right)\quad ?\) what about the \(\bf sin\left(\frac{\pi}{2}\right)\quad ?\)

OpenStudy (anonymous):

Isn't cos 1/2?

OpenStudy (jdoe0001):

what do you mean?

OpenStudy (anonymous):

the cos (pi/2). is it 1/2?

OpenStudy (jdoe0001):

is that what you get from the Unit Circle?

OpenStudy (anonymous):

I'm trying to recall a chart we made in class regarding this. And I believe that it is 1/2

OpenStudy (jdoe0001):

well... you may want to get a Unit Circle :) if you don't have one there are many online btw

OpenStudy (anonymous):

on the unit circle it says pi/2 is 90 degrees and it is located at (0,1)

OpenStudy (jdoe0001):

ok.... so.... the cosine... value will be?

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

http://www.sciencedigest.org/UnitCircle.gif <---- check this unit circle, it has the pair as cosine,sine keep in mind that cosine is just the ADJACENT SIDE's length and sine is the OPPOSITE SIDE's length on the given angle (x,y) => (cosine, sine)

OpenStudy (anonymous):

Oh alright I had it backwards. So cos is 0 and sin is 1. What would be the next step in proving the identity?

OpenStudy (jdoe0001):

ok... so now we know that \(\bf cos\left(\frac{\pi}{2}\right)=0\ and\ sin\left(\frac{\pi}{2}\right)=1\) thus \(\bf \cfrac{cos(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}+sin(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}} {sin(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}-cos(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}\\ \quad \\ \implies \cfrac{cos(x){\color{blue}{0}}+sin(x){\color{blue}{ 1}}} {sin(x){\color{blue}{ 0}}-cos(x){\color{blue}{ 1}}}\) what would you get for that?

OpenStudy (anonymous):

Are we trying to find the tangent? So x/y? or 0/1? Which would just be 0, correct?

OpenStudy (jdoe0001):

verify the identity..... as in make the left-side like the right-side or the other way around

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}+sin(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}} {sin(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}-cos(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}\\ \quad \\ \implies \cfrac{cos(x)\cdot {\color{blue}{0}}+sin(x)\cdot {\color{blue}{ 1}}} {sin(x)\cdot {\color{blue}{ 0}}-cos(x)\cdot {\color{blue}{ 1}}}\)

OpenStudy (jdoe0001):

looks closely

OpenStudy (jdoe0001):

look rather... anyhow... so...what do you think?

OpenStudy (anonymous):

I apologize, you're doing a very good job of explaining this, I'm just still completely lost.

OpenStudy (jdoe0001):

recall that SOMETHING multiplied by 0 is = ? and SOMETHING multiplied by 1 = ?

OpenStudy (anonymous):

so it's 0?

OpenStudy (jdoe0001):

lemme.... rewrite it some

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}+sin(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}} {sin(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}-cos(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}\\ \quad \\ \implies \cfrac{cos(x)\cdot {\color{blue}{0}}+sin(x)\cdot {\color{blue}{ 1}}} {sin(x)\cdot {\color{blue}{ 0}}-cos(x)\cdot {\color{blue}{ 1}}}\qquad cos(x)=a\qquad sin(x)=b\\ \quad \\ \implies \cfrac{(a\cdot 0)-(b\cdot 1)}{(b\cdot 0)-(a\cdot 1)}\) what would you get from that fraction?

OpenStudy (anonymous):

1?

OpenStudy (jdoe0001):

well.... how did you get 1?

OpenStudy (jdoe0001):

hmmm actually it's supposed to be a + above... so \(\bf \cfrac{(a\cdot 0)+(b\cdot 1)}{(b\cdot 0)-(a\cdot 1)}\)

OpenStudy (anonymous):

0 multiplied by anything is 0. So is the answer -1?

OpenStudy (jdoe0001):

-1? is it? whatever happend to the variables?

OpenStudy (jdoe0001):

hehehe, the variables went MIA =)

OpenStudy (anonymous):

\[-\frac{ \sin }{ \cos }\]

OpenStudy (jdoe0001):

ahhhh... notice.... \(\bf tan(\theta)=\cfrac{sin(\theta)}{cos(\theta)}\qquad thus\qquad -\cfrac{sin(\theta)}{cos(\theta)}\implies -tan(\theta)\)

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}+sin(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}} {sin(x){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}-cos(x){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}\\ \quad \\ \implies \cfrac{cos(x)\cdot {\color{blue}{0}}+sin(x)\cdot {\color{blue}{ 1}}} {sin(x)\cdot {\color{blue}{ 0}}-cos(x)\cdot {\color{blue}{ 1}}}\implies -\cfrac{sin(x)}{cos(x)}\implies -tan(x)\)

OpenStudy (anonymous):

Wow okay! So that is how the proof is from start to finish?

OpenStudy (anonymous):

Thank you so much! You're awesome!